OpenStudy (anonymous):

Simplify -3sin(x)-4sin(x)

OpenStudy (tkhunny):

Factor out sin(x), usung the distributive property of multiplication over addition.

OpenStudy (anonymous):

Sorry -3sin(x)-4sin(2x)

OpenStudy (tkhunny):

sin(2x) = 2sin(x)cos(x), then what I said before.

OpenStudy (anonymous):

Thanks will you help me solve ln(8x)-x=0

OpenStudy (tkhunny):

It can't be done conveniently. Are you prepared to utilize numeric algorithms?

OpenStudy (anonymous):

Thank you but no

OpenStudy (anonymous):

But then can you help me find the points of inflection of 3sin(x)+sin(2x) please

OpenStudy (tkhunny):

Is ther a rule about one questin per thread? You'll need a 1st derivative. Let's see what you get.

OpenStudy (anonymous):

Sin(x)(-11cos(x)?

OpenStudy (anonymous):

Hello

OpenStudy (tkhunny):

How did you get that? (d/dx)(3sin(x)+sin(2x)) = 3cos(x) + cos(2x)*2

OpenStudy (anonymous):

Oh i took the derivative of the one you just wrote. I think

OpenStudy (anonymous):

So the second derivative ..

OpenStudy (tkhunny):

(d/dx)(3cos(x) + cos(2x)*2) = -3sin(x) - 2sin(2x)*2 And we have where this conversation started.

OpenStudy (anonymous):

Good work

OpenStudy (anonymous):

So did i get the first deriativ right?

OpenStudy (anonymous):

I mean second sorry agaim

OpenStudy (tkhunny):

Seems like it, but I don't actually know the question, so it's pretty hard to say.

OpenStudy (anonymous):

Do you remember how to find points of inflection?

OpenStudy (anonymous):

Are you old

OpenStudy (anonymous):

Im sorry for the trouble

OpenStudy (anonymous):

If you want you dont hav to help

OpenStudy (anonymous):

Just say goodbye

OpenStudy (anonymous):

:)

OpenStudy (tkhunny):

Oh, are we finding points of inflection? You really shoudl state the entire problem in the beginning. A piece at a time is a little annoying. So, there you have it. -3sin(x) - 2sin(2x)*2 = -3sin(x) - 4sin(2x) = -3sin(x) - 8sin(x)cos(x) = -sin(x)(3 + cos(x)) Don'y spend too much time ont he second factor! Okay, now I'm out.